# Evolutionary Optimization of Physics-Informed Neural Networks: Survey and Prospects

### v2

@@ Abstract
Deep learning models trained on finite data lack a complete understanding of the physical world. On the other hand, physics-informed neural networks (PINNs) are infused with such knowledge through the incorporation of mathematically expressible laws of nature into their training loss function. By complying with physical laws, PINNs provide advantages over purely data-driven models in limited-data regimes. This feature has propelled them to the forefront of scientific machine learning, a domain characterized by scarce and costly data. However, the vision of accurate physics-informed learning comes with significant challenges. This review examines PINNs for the first time in terms of model optimization and generalization, shedding light on the need for new algorithmic advances to overcome issues pertaining to the training speed, precision, and generalizability of today’s PINN models. Of particular interest are the gradient-free methods of neuroevolution for optimizing the uniquely complex loss landscapes arising in PINN training. Methods synergizing gradient descent and neuroevolution for discovering bespoke neural architectures and balancing multiple conflicting terms in physics-informed learning objectives are positioned as important avenues for future research. Yet another exciting track is to cast neuroevolution as a meta-learner of generalizable PINN models.
@@
Physics-informed neural networks, physics-informed neuroevolution, evolutionary optimization, generalizability, survey.
@@ I Introduction
Over the past decade, advances in deep learning have garnered increasing attention for their potential to address pressing challenges in science and engineering. The emerging topic of artificial intelligence (AI) for science promises to transform technological frontiers [1]. For example, learned models have shown the capacity to accelerate the design and discovery of novel materials, photonic devices, proteins, drugs, and renewable energy technologies such as wind turbines. However, unlike conventional applications of AI in vision, speech, or natural language processing where there is a deluge of available data, applications in scientific domains are often characterized by extreme data scarcity. This has raised concerns regarding the veracity of such models when trained only on small data. The cost of acquiring quality data gives rise to a practical bottleneck, leading to naive statistical models which sometimes generate physically inconsistent outputs.
This review focuses on the growing body of literature on physics-informed neural networks (PINNs), popularized in recent years by the work of Raissi et al. [12]. The essential idea of a PINN—where mathematically-defined learning biases are directly incorporated into a neural network’s loss function—is simultaneously decades old, with early work in the 1990s [13, 14], and also a breath of fresh air for AI. This physics-informed loss is amenable to various forms of scientific knowledge and theories, including conservation laws and/or governing equations in the form of ordinary and partial differential equations (ODEs and PDEs), yet retains immense flexibility in the choice of network architectures (e.g., multilayer perceptron, convolutional or graph neural networks) and learning algorithms (e.g., gradient-based or gradient-free). This makes it possible for scientific knowledge uncovered over centuries, from the micro-scale evolution of quantum-mechanical systems to the macro-scale behavior of solids, fluids and propagation of electromagnetic fields, to be incorporated into the learned models of today.
@@
PINN for Fluid Mechanics Turbomachinery Design [15]
@@
PINN for Solid Mechanics Topology Optimization [17, 18]
@@
PINN for Chemistry & Electrochemistry Battery Management Systems [24]
PINN for Quantum & Particle Physics Diffuse Optical Tomography [25]
@@
PINN for Material Science Composite-Tool Systems during Manufacture [28]
@@
Despite the broad applicability of PINNs (ref. to Table I for a detailed summary of potential applications of PINNs), the switch from standard data-driven loss functions to physics-informed learning objectives has brought unforeseen difficulties in optimizing the latter’s uniquely complex landscape [34, 35, 36, 37, 38]. These training difficulties undermine the theoretical advantages of PINNs, and are currently a barrier to their adoption in place of established numerical methods in science and engineering. Existing surveys [8, 39, 40, 41, 42, 43] review PINN methods, applications, benchmark data, limitations, trends and prospects from interdisciplinary viewpoints, but a common point of emphasis is that efficient optimization of PINNs remain an open question direction [43].
@@
In light of the above, this review paper examines PINNs for the first time through the lens of their optimization and generalization challenges. Notably, most PINN implementations today make use of gradient-based optimization algorithms. The ubiquity of the method may be attributed to inertia from the primacy of stochastic gradient descent (SGD) in conventional data-rich deep learning tasks. However, SGD is no silver bullet. Relative to standard deep learning, physics-informed loss functions are not only highly non-convex but also exhibit disparities in the magnitudes of back-propagated gradient signals between the data and physics loss terms (see II-A). These issues lead to pathologies that hinder effective training [44, 45]. The resulting optimization trajectory is littered with spurious minima, making it extremely challenging to find globally optimum solutions that satisfy physics. There is a high chance for gradient descent to get trapped in a local minimum that fails to satisfy the requisite physics, a possibility discussed in greater detail in Section III.
@@
Figure 2: Boosting scientific ML with synergistic physics-informed neuroevolution (PINE) learning.
@@
In order to fulfill the vision of data-efficient scientific machine learning, this paper asserts the need for algorithmic advances beyond gradient descent to move the needle for PINNs. Fig. 2 provides a schematic summary of some of the main computational challenges in this regard, and some corresponding proposed solutions. In Section IV, arguments are presented in support of our main claim: that algorithmic ideas developed within the computational intelligence community may be uniquely well-suited for physics-informed learning. Of particular interest are the gradient-free algorithms of neuroevolution that suffer less from spurious local minima due to their exploration-oriented global search capacity [46, 47]. In addition to optimizing network parameters, neuroevolution enables the discovery of bespoke architectures for physics-informed learning [48]. Evolutionary mechanisms could also prove useful for guiding training data generation in areas with higher errors on governing physics equations, having shown evidence of improved convergence to correct solutions [49].
@@
An associated research direction is the synergy of neuroevolution with gradient-based training (in the spirit of memetic algorithms), marrying the best of both worlds. While population-based neuroevolution excels at navigating highly non-convex loss landscapes and balancing trade-offs between conflicting terms in physics-informed learning objectives, local search via gradient descent can facilitate rapid convergence to precise solutions. Yet another exciting track is to cast neuroevolution as a meta-learner of generalizable PINN models [50]. This may be achieved through a bilevel formulation where an upper-level evolutionary algorithm searches for PINN configurations that can be quickly fine-tuned for various downstream physics tasks at the lower-level.
@@
The advancement of physics-informed neuroevolution is expected to have far-reaching impacts on establishing PINN methodologies at the forefront of scientific machine learning. Yet, surprisingly, there are only a handful of related works in the literature that explore such ideas, implying huge scope for future research and development.
@@ II-A Physics-Informed Loss
$\mathcal{L}_{PINN}=\lambda_{physics}\ \mathcal{L}_{physics}+\lambda_{data}\ \mathcal{L}_{data},$ (2)
@@ II-B Physics-Informed Learning Advantages
The advantages of physics-informed learning are demonstrated in Fig. 4. Given only a few labeled data ($n$=$5$), a DNN model (with sufficient representation capacity) remains unable to correctly reconstruct the entire system, although it fits very well on all training samples. On the other hand, a PINN model can accurately learn the solution from the same set of labeled data. Given that the physics constraints are enforced on a set of label-free training samples ($n_{pde}$=$1000$) uniformly drawn from the computation domain utilizing the aforementioned PINN loss, a PINN can learn the solution even without any labeled data. The governing physics in this example is the convection-diffusion equation $\alpha\frac{du}{dx}-\frac{d^{2}u}{dx^{2}}=0$ with PDE parameter $\alpha$, which is a ubiquitous physics model in science and engineering that describes the distribution of a scalar quantity (e.g., mass, energy, temperature) in the presence of convective transport and diffusion [53].
@@ II-C PINN for Scientific Simulation
While much of the PINN literature focuses on single problems, a PINN model can also incorporate additional parameterizations in the input space, allowing for learning from and solution across a parametric family of physics scenarios [56, 57, 58, 59]. These parameterizations can include geometrical aspects or physically meaningful system parameters, such as the Reynolds number or material dielectric constant. These models, also referred to as parameterized PINN, can then provide predictions across a set of physics scenarios as opposed to providing the solution for a single scenario.
@@ II-D PINN for Scientific Discovery
The PINN models can be applied to:
  1. 1.
    Inference of system-level parameters such as the underlying governing equations or constitutive relationships or system parameters (e.g., Reynolds number, wave speed) based on observations [60, 61, 62, 63, 64, 65].
  2. 2.
    Inference of unobserved field variables across the entire domain based on limited observations, e.g., predicting pressure in a particle imaging velocimetry experiment with only velocity observations [66]. Many have employed PINNs towards the inference of domain field variables in various problems spanning fluid dynamics [67], heat transfer [68] and biomedical sciences [69, 70].
  3. 3.
    PDE-constrained design optimization towards a set of desired (observation) outcomes. The PINN is used as an adjoint optimizer, enabling the design parameters to be optimized towards the desired outcome while fulfilling all specified relevant physics during training. This has been successfully demonstrated on topology optimization [71], metamaterial [72] and biomimetics [73] design problems.
PINN models are uniquely suited for hypothesis inference when compared to numerical solvers and can effectively synergize theory and observation data from various sources such as sensors and imaging, even for ill-posed PDE problems. It can be combined with probabilistic inference methods (e.g. Bayesian-PINN [74]), with the probabilistic framework offering a mathematical tool to handle uncertainties arising from both noisy data and physics assumptions [74, 75, 76].
@@ III Optimization and Generalization Challenges
III-A Challenges in Optimizing PINN Model
@@ III-A1 Complex Loss Landscape
In addition, the PINN optimization landscapes are littered with spurious minima and are more complex than those of usual data-driven DNN models. Fig. 5 illustrates the loss landscapes of PINN and DNN models for the convection-diffusion problem presented in Subsection II-B. This visualization provides evidence of a much higher degree of complexity and ruggedness of the loss landscape for PINN compared to usual data-driven DNN models. The smoothness of the latter is conducive to gradient-based algorithms that can easily find descent paths from the point of initialization to a good local optimum that fits the training data well. This outcome corroborates the primacy of SGD and its variants in today’s deep learning tasks. In contrast, the contradictory descent directions from different loss terms encountered during PINN training can adversely impact optimization [80, 81].
Although these powerful gradient-based algorithms are inseparable from the recent success of deep learning, they are no silver bullet. Gradient-based methods are susceptible to being trapped in spurious local minima, which is often the case in the context of PINN optimization. Moreover, usual deep learning applications do not require a fully converged solution because they prioritize generalization performance over training precision. However, this is not the case for PINN, where the loss needs to be as close to the global minimum as possible to ensure the veracity of the model concerning the underlying physics of a problem. A small violation of the PDE constraints can result in an unphysical PINN solution while a small violation of the BC or IC constraints corresponds to an entirely different PDE problem [82, 83]. This issue is especially detrimental in time-varying dynamical problems where small violations early in time can cause errors to propagate and accumulate in later times.
@@ III-A2 Initial Local Minimum Trap and Spectral Bias
PINN models using common weights initialization (such as Xavier initialization) and activation function (such as tanh, sigmoid) are initialized around flat output functions, and are consequently susceptible to a local minimum trap at initialization whereby the PINN loss already sees a minimized PDE loss term; see theoretical analysis in [84, 85]. Such a strong bias around initial flat functions can be difficult for gradient descent optimization to escape, especially when sensible solutions that simultaneously minimize all the PDE and IC/BC loss terms are not accessible near the point of initialization.
Another closely related issue stems from the known saturation property of the activation functions commonly used for PINN, such as tanh and sigmoid. Whenever the network weights are shifted towards a larger distribution, e.g., as driven by the need to approximate higher frequencies and steep gradients common in physics, such saturation property can cause their backpropagation gradients to vanish. This makes further optimization of the network weights hard for gradient descent, causing a learning bias towards smooth functions. This known tendency—also called spectral bias—has been studied through the lens of NTK [86, 87], revealing a significant discrepancy in convergence rates for different frequency components in the output during training (lower frequencies are learned before higher frequencies). While this has been regarded as a natural regularizer with potential benefits to the generalization performance of DNNs, the consequences are highly deleterious for PINNs. Spectral bias makes it difficult for PINNs to accurately learn complex nonlinear physics behaviors [88], e.g., in the vicinity of steep gradients characterizing shock waves. These local discrepancies further deter propagation of the correct initial and boundary conditions into the computation domain, which has been suggested to be critical for successful PINN training [83, 89], and eventually becomes a bottleneck to convergence of the PINN loss.
This saturation also translates to increased difficulties for gradient-based algorithms to alter the behavior of PINN output far away from the input origin. As a result, PINN struggles to learn over long time horizons, partly because the learning bias towards smooth functions can cause the gradient descent to prematurely converge to a trivial solution (satisfying the PDE) before the initial and boundary conditions can properly propagate to a later time [90].

### v3

# Evolutionary Optimization of Physics-Informed Neural Networks: Evo-PINN Frontiers and Opportunities

@@ Abstract
Deep learning models trained on finite data lack a complete understanding of the physical world. On the other hand, physics-informed neural networks (PINNs) are infused with such knowledge through the incorporation of mathematically expressible laws of nature into their training loss function. By complying with physical laws, PINNs provide advantages over purely data-driven models in limited-data regimes and present as a promising route towards Physical AI. This feature has propelled them to the forefront of scientific machine learning, a domain characterized by scarce and costly data. However, the vision of accurate physics-informed learning comes with significant challenges. This work examines PINNs for the first time in terms of model optimization and generalization, shedding light on the need for new algorithmic advances to overcome issues pertaining to the training speed, precision, and generalizability of today’s PINN models. Of particular interest are gradient-free evolutionary algorithms (EAs) for optimizing the uniquely complex loss landscapes arising in PINN training. Methods synergizing gradient descent and EAs for discovering bespoke neural architectures and balancing multiple terms in physics-informed learning objectives are positioned as important avenues for future research. Another exciting track is to cast evolutionary as a meta-learner of generalizable PINN models. To substantiate these proposed avenues, we further highlight results from recent literature to showcase the early success of such approaches in addressing the aforementioned challenges in PINN optimization and generalization.
@@
Physics-informed neural networks, evolutionary optimization, neuroevolution, generalizability, survey.
@@ I Introduction
Over the past decade, advances in deep learning have garnered increasing attention for their potential to address pressing challenges in science and engineering. The emerging topic of artificial intelligence (AI) for science promises to transform technological frontiers [1]. For example, learned models have shown the capacity to accelerate the design and discovery of novel materials, photonic devices, proteins, drugs, and renewable energy technologies such as wind turbines. However, unlike conventional applications of AI in vision, speech, or natural language processing where there is a deluge of available data, applications in scientific domains are often characterized by extreme data scarcity. This has raised concerns regarding the veracity of such models when trained only on small data. The cost of acquiring quality data gives rise to a practical bottleneck, leading to naive statistical models which sometimes generate physically inconsistent outputs. These shortcomings become even more consequential in the context of Agentic AI systems [2, 3, 4], which are designed to autonomously drive scientific discovery and decision-making. In such systems, a lack of embedded physical understanding by models can lead to flawed reasoning or unsafe actions when interfacing with the real world [5, 6].
@@
Fluid Mechanics Turbomachinery Design [7]
@@
Solid Mechanics Topology Optimization [9, 10]
@@
Chemistry & Electrochemistry Battery Management Systems [16]
Quantum & Particle Physics Diffuse Optical Tomography [17]
@@
Material Science Composite-Tool Systems during Manufacture [20]
This effort focuses on the growing body of literature on physics-informed neural networks (PINNs), popularized in recent years by the work of Raissi et al. [31]. The essential idea of a PINN—where mathematically-defined learning biases are directly incorporated into a neural network’s loss function—is simultaneously decades old, with early work in the 1990s [32, 33], and also a breath of fresh air for AI. This physics-informed loss is amenable to various forms of scientific knowledge and theories, including conservation laws and/or governing equations in the form of ordinary and partial differential equations (ODEs and PDEs), yet retains immense flexibility in the choice of network architectures (e.g., multilayer perceptron, convolutional or graph neural networks) and learning algorithms (e.g., gradient-based or gradient-free). This makes it possible for scientific knowledge uncovered over centuries, from the micro-scale evolution of quantum-mechanical systems to the macro-scale behavior of solids, fluids and propagation of electromagnetic fields, to be incorporated into the learned models of today.
@@
Despite the broad applicability of PINNs (ref. to Table I for a summary of potential applications of PINNs), the switch from standard data-driven loss functions to physics-informed learning objectives has brought unforeseen difficulties in optimizing the latter’s uniquely complex landscape [39, 40, 41, 42, 43]. These training difficulties undermine the theoretical advantages of PINNs, and are currently a barrier to their adoption in place of established numerical methods in science and engineering. Existing surveys [27, 44, 45, 46, 47, 48] review PINN methods, applications, benchmark data, limitations, trends and prospects from interdisciplinary viewpoints, but a common point of emphasis is that efficient optimization of PINNs remain an open question direction [48].
@@
Figure 2: Boosting scientific ML with evolutionary optimization of physics-informed neural networks (Evo-PINN).
In light of the above, this paper examines PINNs for the first time through the lens of their optimization and generalization challenges. Notably, most PINN implementations today make use of gradient-based optimization algorithms. The ubiquity of the method may be attributed to inertia from the primacy of stochastic gradient descent (SGD) in conventional data-rich deep learning tasks. However, SGD is no silver bullet. Relative to standard deep learning, physics-informed loss functions are not only highly non-convex but also exhibit disparities in the magnitudes of back-propagated gradient signals between the data and physics loss terms (see II-A). These issues lead to pathologies that hinder effective training [49, 50]. The resulting optimization trajectory is littered with spurious minima, making it extremely challenging to find globally optimum solutions that satisfy physics. There is a high chance for gradient descent to get trapped in a local minimum that fails to satisfy the requisite physics, a possibility discussed in greater detail in Section III.
@@
In order to fulfill the vision of data-efficient scientific machine learning, this paper asserts the need for algorithmic advances beyond gradient descent to move the needle for PINNs. Fig. 2 provides a schematic summary of some of the main computational challenges in this regard, and some corresponding proposed solutions. In Section IV, arguments are presented in support of our main claim: that algorithmic ideas developed within the computational intelligence community may be uniquely well-suited for physics-informed learning. Of particular interest are the gradient-free evolutionary algorithms (EAs) that suffer less from spurious local minima due to their exploration-oriented global search capacity [51, 52]. In addition to optimizing network parameters, EAs enable the discovery of bespoke architectures for physics-informed learning [53]. Evolutionary mechanisms could also prove useful for guiding training data generation in areas with higher errors on governing physics equations, having shown evidence of improved convergence to correct solutions [54].
@@
An associated research direction is the synergy of EAs with gradient-based training (in the spirit of memetic algorithms), thereby marrying the best of both worlds. While population-based EAs may be advantageous in navigating highly non-convex loss landscapes and balancing trade-offs between conflicting terms in physics-informed learning objectives, local search via gradient descent can facilitate rapid convergence to precise solutions. Yet another exciting track is to cast evolutionary as a meta-learner of generalizable PINN models [55]. This may be achieved through a bilevel formulation where an upper-level evolutionary searches for PINN configurations that can be quickly fine-tuned for various downstream physics tasks at the lower-level.
@@
The advancement of Evo-PINNs is expected to have far-reaching impacts on establishing PINN methodologies at the forefront of scientific machine learning. Yet, surprisingly, there are only a handful of related works in the literature that explore such ideas, implying huge scope for future research and development.
@@ II-A Physics-Informed Loss
$\mathcal{L}_{\text{PINN}}=\lambda_{physics}\ \mathcal{L}_{physics}+\lambda_{data}\ \mathcal{L}_{data},$ (2)
@@ II-B Physics-Informed Learning Advantages
The advantages of physics-informed learning are demonstrated in Fig. 4. Given only a few labeled data ($n$=$5$), a DNN model (with sufficient representation capacity) remains unable to correctly reconstruct the entire system, although it fits very well on all training samples. On the other hand, a PINN model can accurately learn the solution from the same set of labeled data. Given that the physics constraints are enforced on a set of label-free training samples ($n_{pde}$=$1000$) uniformly drawn from the computation domain utilizing the aforementioned PINN loss, a PINN can learn the solution even without any labeled data. The governing physics in this example is the convection-diffusion equation $\alpha\frac{du}{dx}-\frac{d^{2}u}{dx^{2}}=0$ with PDE parameter $\alpha$, a ubiquitous physics model in science and engineering that describes the distribution of a scalar quantity (e.g., mass, energy, temperature) in the presence of convective transport and diffusion [58].
@@ II-C PINN for Scientific Simulation
While most PINN studies address single physics problem, a PINN model can also incorporate additional input parameters, such as geometry, Reynolds number, or material properties [61, 62, 63, 64]. These models, also referred to as parameterized PINN, can then provide predictions across a set of physics scenarios, rather than just one.
@@ II-D PINN for Scientific Discovery
The PINN models are uniquely suited for hypothesis inference when compared to numerical solvers and can effectively synergize theory and observation data from various sources such as sensors and imaging, even for ill-posed PDE problems. It can be combined with probabilistic inference methods (e.g. Bayesian-PINN [65]), with the probabilistic framework offering a mathematical tool to handle uncertainties arising from both noisy data and physics assumptions [65, 66, 67].
In the literature, PINN models have been applied to: (1) Inference of system-level parameters such as the underlying governing equations or constitutive relationships or system parameters (e.g., Reynolds number, wave speed) based on observations [68, 69, 70, 71, 72, 73]. (2) Inference of unobserved field variables across the entire domain based on limited observations, e.g., predicting pressure in a particle imaging velocimetry experiment with only velocity observations [74]. Many have employed PINNs towards the inference of domain field variables in various problems spanning fluid dynamics [75], heat transfer [76] and biomedical sciences [77, 78]. (3) PDE-constrained design optimization towards a set of desired (observation) outcomes. The PINN is used as an adjoint optimizer, enabling the design parameters to be optimized towards the desired outcome while fulfilling all specified relevant physics during training. This has been successfully demonstrated on topology optimization [79], metamaterial [80] and biomimetics [81] design problems.
@@ III Optimization and Generalization Challenges
III-A Challenges in Optimizing PINN Models
@@ III-A1 Complex Loss Landscape
In addition, the PINN optimization landscapes are littered with spurious minima and are more complex than those of usual data-driven DNN models. Fig. 5 illustrates the loss landscapes of PINN and DNN models for the convection-diffusion problem presented in Subsection II-B. To investigate the topological differences in their loss landscapes, we first computed the Hessians of their respective loss functions (i.e., $\mathcal{L}_{\text{PINN}}$ or $\mathcal{L}_{\text{DNN}}$) with respect to the model parameters: $\mathbf{H}=\nabla^{2}\ \mathcal{L}(\boldsymbol{\boldsymbol{w}})$. We then extracted the leading two eigenvectors and eigenvalues of the Hessian (i.e., first 2 principal Hessian directions) which correspond to the directions of highest curvature in the loss landscape. By perturbing the trained models along these dominant eigenvectors and evaluating the resulting loss values, we can obtain a more informative characterization of the local topology than perturbing the model parameters in random directions [39]. This visualization provides evidence of a much higher degree of complexity and ruggedness of the loss landscape for PINN compared to usual data-driven DNN models. The smoothness of the latter is conducive to gradient-based algorithms that can easily find descent paths from the point of initialization to a good local optimum that fits the training data well. This outcome corroborates the primacy of SGD and its variants in today’s deep learning tasks. In contrast, the contradictory descent directions from different loss terms encountered during PINN training can adversely impact optimization [85, 86]. It is worth nothing that when the loss terms are purely physics-based (e.g., PDE and IC/BC), they are not theoretically contradictory. However, during the optimization process, especially when the solution is still far from the true optimum, the local gradient directions of these loss components can still become practically conflicting. This issue arises due to the combined effects of neural network expressivity, the structure of the loss landscape (e.g., scales and convergence dynamics), and the limitations of optimizers, which can cause certain objectives to dominate or suppress others during early or intermediate stages of training.
Although these powerful gradient-based algorithms are inseparable from the recent success of deep learning, they are no silver bullet. For instance, gradient-based methods are susceptible to being trapped in spurious local minima, which is often the case in the context of PINN optimization. Typical deep learning applications do not require a fully converged solution, as they prioritize generalization performance over training precision. However, this is not the case for PINNs, where reaching the global minimum is essential to ensure the model’s fidelity to the underlying physics. A small violation of the PDE constraints can result in an unphysical PINN solution while a small violation of the BC or IC constraints corresponds to an entirely different PDE problem [87, 88].
@@ III-A2 Initial Local Minimum Trap and Spectral Bias
PINN models employing standard activation functions (such as tanh, sigmoid) when initialized using common techniques (such as the Xavier method) tend to produce flat output functions at initialization, making them immediately susceptible to a local minimum trap whereby the PINN loss already sees a minimized PDE loss term. This theoretical insight, discussed in [89, 90, 91], applies across a range of physics scenarios where the underlying PDE is trivially satisfied by flat functions. A strong bias around flat functions can therefore be difficult for gradient descent to escape, especially when the true solution, that must simultaneously minimize all PDE and IC/BC loss terms, is not accessible near the point of initialization.
Another closely related issue stems from the known saturation property of the activation functions commonly used for PINN. Such saturation property can cause their backpropagation gradients to vanish whenever the network weights are shifted towards a larger distribution, as driven by the need to approximate higher frequencies and steep gradients common in physics. This makes further gradient descent of the network weights difficult, adversely causing a learning bias towards smooth functions (lower frequencies are learned better than higher frequencies). This known tendency—also called spectral bias—has been studied through the lens of NTK [92, 93], revealing a significant discrepancy in convergence rates for different frequency components in the output during training. While this has been regarded as a natural regularizer with potential benefits to the generalization performance of DNNs, the consequences are highly deleterious for PINNs. Spectral bias makes it difficult for PINNs to accurately learn complex nonlinear physics behaviors [94], e.g., in the vicinity of steep gradients characterizing shock waves. These local discrepancies further deter propagation of the correct initial and boundary conditions into the computation domain, which has been suggested to be critical for successful PINN training [88, 95].
This saturation also translates to increased difficulties for gradient-based algorithms to alter the behavior of PINN output far away from the input origin. As a result, PINN struggles to learn over long time horizons, partly because the learning bias towards smooth functions can cause the gradient descent to prematurely converge to a trivial solution (satisfying the PDE) before the initial and boundary conditions can properly propagate to a later time [96]. An associated tendency for gradient-based PINN training to be drawn towards unstable fixed point solutions of dynamical systems was analyzed in [97], with experiments suggesting that they form strong basins of attraction in the PINN loss landscape.